a chemical compound has 54.55% carbon and 9.09% hydrogen. Determine the chemical formula if its molar mass is 70g / mol
1
Expert's answer
2020-09-21T06:49:31-0400

Solution.

"n(C) = \frac{54.55}{12} = 4.546"

"n(H) = \frac{9.09}{1} = 9.09"

"n(C):n(H):n(X) = 3.18:6.36: \frac{25.452}{A(X)} = 1:2:\frac{8}{A(X)}"

By iterating from the element nitrogen to chlorine, the most suitable formula for this compound is C3H6N2. So, X is N.

Answer:

C3H6N2